OpenStudy (gabbyalicorn):

A circle has a radius of 12 m. What is the approximate area of the circle? Use 3.14 to approximate pi. Write the answer to the nearest hundredth.

OpenStudy (gabbyalicorn):

so 3.14 x 12 x 2 = 75.36 / 2 = 37.68

OpenStudy (gabbyalicorn):

is tht correct?

OpenStudy (gabbyalicorn):

@10115658

OpenStudy (rainbow_rocks03):

area of a circle is pir^2 A=π(r)2 where A is area, pi is 3.14 and r is radius

OpenStudy (rainbow_rocks03):

okay so try to work it out, and show me wh at you get

OpenStudy (rainbow_rocks03):

and you got 3.14 x 12 x 2 = 75.36 / 2 = 37.68

OpenStudy (gabbyalicorn):

i got 6.28 now

OpenStudy (rainbow_rocks03):

no no no don't do that you were on track

OpenStudy (gabbyalicorn):

37.68?

OpenStudy (rainbow_rocks03):

so you multiplied by 12, to get 37.68 multiply your answer by 12 again

OpenStudy (gabbyalicorn):

452.16

OpenStudy (rainbow_rocks03):

correct ! I also copied and pasted it for us to understand

OpenStudy (gabbyalicorn):

Thanks for your help!

OpenStudy (gabbyalicorn):

wait wait wai!

OpenStudy (rainbow_rocks03):

wait what?

OpenStudy (gabbyalicorn):

it says approximate does that mean we round

OpenStudy (rainbow_rocks03):

IDK

OpenStudy (gabbyalicorn):

omg i might get it wrong then :(

OpenStudy (rainbow_rocks03):

I am 89% that is the answer but the answer might either be 452.16 or 37.68

OpenStudy (rainbow_rocks03):

BUT YOU WILL PASS I BEILIEVE IN YOU!

OpenStudy (gabbyalicorn):

0.0 can u tag some1 else to double check pls im scared now XD

OpenStudy (gabbyalicorn):

it says "Write the answer to the nearest hundredth." please help me!

OpenStudy (gabbyalicorn):

@rainbow_rocks03 @10115658